# Quadratic Programming-based Reference Spreading Control for Dual-Arm Robotic Manipulation with Planned Simultaneous Impacts

### v3

@@ I-B Related work on impact-aware tracking control
While RS control was initially defined for single impact events, an approach to deal with nominally simultaneous impacts has been introduced in [20]. Uncertainty in the environment could cause a sequence of impacts rather than the planned nominally simultaneous impact, resulting in unexpected velocity jumps and a contact state that neither corresponds to the ante-impact nor the post-impact state. To avoid peaks and steps in the actuator inputs that would occur when naively applying velocity feedback, an additional interim-impact mode was defined. This interim mode is active in the time frame when contact between two impacting bodies has only been partially established, and uses a feedforward signal based on the ante-impact mode without any position or velocity feedback to be independent from any jumps in velocity. In [15], an extension to the RS framework was proposed by defining a novel interim mode, which also uses position feedback during this interim mode to pursue persistent contact establishment without relying on velocity feedback. Furthermore, the approach in [15] is cast into the quadratic programming (QP) robot control framework [28, 29, 30], where the control input is obtained from a linearly constrained quadratic optimization problem. QP control can be used for control of robots with multiple tasks in joint space or operational space, under a set of constraints ensuring, for example, adherence to joint limits or avoidance of unwanted collisions. This is essential in real-life robotic applications, and allows the control approach to be applied on complex robotic systems, such as humanoids or quadrupeds. A control algorithm based on RS using time-invariant references rather than the traditionally used time-based references has been developed in [31, 32], also using a QP control framework. While experimental validation for RS has been provided for a 1 degree of freedom (DOF) mechanical system with single impact events in [33], RS for systems with simultaneous impacts, or for nDOF robots with $n>1$ in general, has not yet been realised.
@@ III Extended reference generation via teleoperation
At the core of RS lies the formulation of ante-impact and post-impact references that are consistent with the impact dynamics and overlap around the impact time. In this section, first, a methodology for obtaining ante- and post-impact references that match the impact dynamics is described, which uses teleoperation with a QP-based impedance controller. Subsequently, a methodology for extending the obtained reference around the nominal impact time is described. This allows for a formulation of the tracking error that does not contain impact-induced peaks or jumps which would translate to the input torques, as mentioned in Section I.
@@ IV Reference spreading based QP control approach
In Section III, a method to acquire a reference motion based on QP-based robot control and teleoperation is detailed together with a concrete example, that is provided for illustration. This motion is used to generate extended ante- and post-impact references that will now be used to construct a RS-based control action without a human in the loop. In this section, we detail how ante-impact, interim, and post-impact control actions can be realized via suitable QPs, obtaining for each mode a suitable joint torque $\bm{\tau}_{i}^{*}$, with $i\in\{1,2\}$, to be applied to the robot arms after every time step $\Delta t$.
@@ V Experimental validation
In order to experimentally validate [^3] A video showcasing the experimental validation described in this section is added as a multimedia attachment. the proposed framework, the control approach described in Section IV is used to autonomously perform a grabbing task for a variety of objects using multiple references generated through the approach described in Section III-B. We have used the QP solver provided by the software framework mc_rtc444mc_rtc: https://jrl-umi3218.github.io/mc_rtc/ to perform these experiments. To emulate uncertainty in the environment, the to-be-manipulated object is shifted in the y-direction of the world frame defined in Figure 1 over a range of distances during different experiments. As depicted in Figure 7, which shows snapshots of an experiment where the object displacement was -30 mm, this can result in a non-simultaneous impact, where one of the robots impacts the object before the other one, and before the nominal impact time $T_{r}$.
@@
Figure 8: Velocity and desired wrench in $y$-direction of the end effector of robot 1 during a representative experiment, compared against the ante- and post-impact velocity reference and the ante- and post-impact feedforward wrench. Black dotted lines indicate mode transitions, with nominal impact time $T_{r}$, recorded impact time $T_{\text{imp}}$, and interim mode duration $\Delta t_{\text{int}}$.
@@
To further validate of the approach, a total of ten different references have been generated to grasp the 1.25 kg rice box using the teleoperation-based approach described in Section III, resulting in a natural variety of grasping motions. Initial experiments using the proposed approach and the three baseline approaches have then been performed for each reference, with a y-displacement of -30 mm and 30 mm, resulting in 80 total experiments. As a result of this displacement, not every experiment results in a successful grasp. Vibrations during contact establishment can result in the object being dropped, or safety limits of the Franka Emika robot [^5] For more information on the safety limits of the Franka Emika robot, see: https://frankaemika.github.io/docs/control_parameters.html#limits-for-panda being exceeded. The success rates for each control approach are depicted in the first column of Table II, where a successful grasp is defined as a grasp where the robot establishes and maintains sustained contact without triggering any safety limit before the prescribed release of the object.
@@
Out of the ten recorded references using the 1.25 kg rice box, three were selected that consistently produce successful experiments for all of the control approaches. On top of this, a successful reference was recorded for three different grasping scenarios using 1) a single 0.64 kg parcel, 2) two parcels with a combined weight of 1.24 kg, and 3) a pack of plastic drink bottles weighing 2.12 kg. All of the test objects are depicted in Figure 9. Experiments with each of the six references were then performed for five different initial object positions, with a y-displacement of -30 mm, -15 mm, 0 mm, 15 mm and 30 mm compared to the object position used during teleoperation. Five experiments have been performed for the proposed control approach and the three baseline approaches for each object displacement value, resulting in a total of $6\times 5\times 5\times 4=600$ additional experiments. The main goal of these experiments is to evaluate the effectiveness of the approach in reducing the peaks and steps in the input forces.
@@
Figure 10: Desired force norm for the proposed control approach and baseline approaches, for an object displacement of -30 mm.
Figure 11: Average desired force norm around the nominal impact time for a total of 600 experiments with different initial object displacements, shown for the proposed control approach and the three baseline approaches.

### v4

@@ I-B Related work on impact-aware tracking control
While RS control was initially defined for single impact events, an approach to deal with nominally simultaneous impacts has been introduced in [20]. Uncertainty in the environment could cause a sequence of impacts rather than the planned nominally simultaneous impact, resulting in unexpected velocity jumps and a contact state that neither corresponds to the ante-impact nor the post-impact state. To avoid peaks and steps in the actuator inputs that would occur when naively applying velocity feedback, an additional interim-impact mode was defined. This interim mode is active in the time frame when contact between two impacting bodies has only been partially established, and uses a feedforward signal based on the ante-impact mode without any position or velocity feedback to be independent from any jumps in velocity. In [15], an extension to the RS framework was proposed by defining a novel interim mode, which also uses position feedback during this interim mode to pursue persistent contact establishment without relying on velocity feedback. Furthermore, the approach in [15] is cast into the quadratic programming (QP) robot control framework [28, 29, 30], where the control input is obtained from a linearly constrained quadratic optimization problem. QP control can be used for control of robots with multiple tasks in joint space or operational space, under a set of constraints ensuring, for example, adherence to joint limits or avoidance of unwanted collisions. This is essential in real-life robotic applications, and is required to eventually extend the RS framework to other more complex robotic systems, such as humanoids or quadrupeds. A control algorithm based on RS using time-invariant references rather than the traditionally used time-based references has been developed in [31, 32], also using a QP control framework. While experimental validation for RS has been provided for a 1 degree of freedom (DOF) mechanical system with single impact events in [33], RS for systems with simultaneous impacts, or for nDOF robots with $n>1$ in general, has not yet been realised.
@@ III Extended reference generation via teleoperation
At the core of RS lies the formulation of ante-impact and post-impact references that are consistent with the impact dynamics and overlap around the impact time. In this section, first, a methodology for obtaining ante- and post-impact references that match the impact dynamics is described, which uses teleoperation with a QP-based impedance controller. We reiterate that the proposed RS control approach is not tied to this teleoperation approach, and can be replaced by an autonomous impact-aware motion planner. Subsequently, a methodology for extending the obtained reference around the nominal impact time is described. This allows for a formulation of the tracking error that does not contain impact-induced peaks or jumps which would translate to the input torques, as mentioned in Section I.
@@ IV Reference spreading based QP control approach
In Section III, a method to acquire a reference motion based on QP-based robot control and teleoperation is detailed together with a concrete example, that is provided for illustration. This motion is then used to generate extended ante- and post-impact references that match the impact dynamics of the system, which will now be used to construct a RS-based control action without a human in the loop. In this section, we detail how ante-impact, interim, and post-impact control actions can be realized via suitable QPs, obtaining for each mode a suitable joint torque $\bm{\tau}_{i}^{*}$, with $i\in\{1,2\}$, to be applied to the robot arms after every time step $\Delta t$.
@@ V Experimental validation
In order to experimentally validate [^3] A video showcasing the experimental validation can be found here: https://youtu.be/jKzKd-z-EGE. the proposed framework, the control approach described in Section IV is used to autonomously perform a grabbing task for a variety of objects using multiple references generated through the approach described in Section III-B. We have used the QP solver provided by the software framework mc_rtc444mc_rtc: https://jrl-umi3218.github.io/mc_rtc/ to perform these experiments. To emulate uncertainty in the environment, the to-be-manipulated object is shifted in the y-direction of the world frame defined in Figure 1 over a range of distances during different experiments. As depicted in Figure 7, which shows snapshots of an experiment where the object displacement was -30 mm, this can result in a non-simultaneous impact, where one of the robots impacts the object before the other one, and before the nominal impact time $T_{r}$.
@@
Figure 8: Velocity and desired wrench in $y$-direction of the end effector of robot 1 during a representative experiment with the rice box object, compared against the ante- and post-impact velocity reference and the ante- and post-impact feedforward wrench. Black dotted lines indicate mode transitions, with nominal impact time $T_{r}$, recorded impact time $T_{\text{imp}}$, and interim mode duration $\Delta t_{\text{int}}$.
@@
To further validate the approach, a total of ten different references have been generated to grasp the 1.25 kg rice box using the teleoperation-based approach described in Section III, resulting in a natural variety of grasping motions. Initial experiments using the proposed approach and the three baseline approaches have then been performed for each reference, with a y-displacement of -30 mm and 30 mm, resulting in 80 total experiments. As a result of this displacement, not every experiment results in a successful grasp. Vibrations during contact establishment can result in the object being dropped, or safety limits of the Franka Emika robot [^5] For more information on the safety limits of the Franka Emika robot, see: https://frankaemika.github.io/docs/control_parameters.html#limits-for-panda being exceeded. The success rates for each control approach are depicted in the first column of Table II, where a successful grasp is defined as a grasp where the robot establishes and maintains sustained contact without triggering any safety limit before the prescribed release of the object.
@@
Out of the ten recorded references using the 1.25 kg rice box, three were selected that consistently produce successful experiments for all of the control approaches. On top of this, a single successful reference was recorded for three different grasping scenarios using 1) a single 0.64 kg parcel, 2) two parcels with a combined weight of 1.24 kg, and 3) a pack of plastic drink bottles weighing 2.12 kg. All of the test objects are depicted in Figure 9. Experiments with each of the six references were then performed for five different initial object positions, with a y-displacement of -30 mm, -15 mm, 0 mm, 15 mm and 30 mm compared to the object position used during teleoperation. Five experiments have been performed for the proposed control approach and the three baseline approaches for each object displacement value, resulting in a total of $6\times 5\times 5\times 4=600$ additional experiments. The main goal of these experiments is to evaluate the effectiveness of the approach in reducing the peaks and steps in the input forces.
@@
Figure 10: Desired force norm for the proposed control approach and baseline approaches, using a rice box with initial displacement of -30 mm.
Figure 11: Average desired force norm around the nominal impact time for a total of 600 experiments with different objects and different initial displacements for each object, shown for the proposed control approach and the three baseline approaches.